OpenStudy (jessiegonzales):

METAL!!!!!!! What is the surface area of a conical grain storage tank that has a height of 37 meters and a diameter of 16 meters? Round the answer to the nearest square meter. A. 2,831 square meters B. 2,664 square meters C. 1,152 square meters D. 1,131 square meters @jim_thompson5910

OpenStudy (anonymous):

Refer to this equation to find the surface area of a cone

OpenStudy (jessiegonzales):

height is 37 and diameter is 8

OpenStudy (anonymous):

The diameter is 16, what is the radius?

OpenStudy (jessiegonzales):

i mean the radius is 8

OpenStudy (anonymous):

\[SA = \pi r^{2} + \pi r l\] You will need to solve for l, the lateral height. r, h, and l make a special shape. How can you use the relationship between r, h, and l to solve for l?

OpenStudy (jessiegonzales):

okay so I multiply 37x8?

OpenStudy (anonymous):

Not quite; what shape do r, h, and l form?

OpenStudy (jessiegonzales):

cone

OpenStudy (anonymous):

On a 3-dimensional perspective, yes absolutely. However, h, r, and l form a right triangle looking from the 2D perspective. You can use the Pythagorean Theorem to solve for l given r and h.

OpenStudy (jessiegonzales):

oh ok

OpenStudy (anonymous):

Solve for l :) Then use the Surface Area equation

OpenStudy (anonymous):

Insight: A cone is basically formed by rotating the triangle 360 degrees, hence why pi is involved.

OpenStudy (jessiegonzales):

ok so how do i solve for l?

OpenStudy (anonymous):

Pythagorean Theorem, your two legs will be h and r. Your hypotenuse will be l. \[a^{2} + b^{2} = c^{2}\] so \[r^{2} + h^{2} = l^{2}\] You already have values for r and h, so solve for l.

OpenStudy (jessiegonzales):

8^2+37^2=L IS THAT RIGHT

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

well acttually no

OpenStudy (anonymous):

You need to square L as well

OpenStudy (jessiegonzales):

64+1369=L^2

OpenStudy (anonymous):

yes, so take the square root of both sides to get L

OpenStudy (anonymous):

\[64 + 1369 = L^{2}\] \[1433 = L^{2}\] \[L = \sqrt{1433}\]

OpenStudy (jessiegonzales):

ok so l=1433

OpenStudy (anonymous):

L is the square root of 1433

OpenStudy (jessiegonzales):

37.8

OpenStudy (anonymous):

This is what you have now: \[r = 8, h = 37, l = \sqrt{1433}\]

OpenStudy (anonymous):

right. So now you use the surface area equation

OpenStudy (anonymous):

\[SA = \pi r^{2} + \pi r l\]

OpenStudy (jessiegonzales):

25.12 for the first part

OpenStudy (anonymous):

Try to leave everything in terms of pi in your calculator, and approximate at the end

OpenStudy (jessiegonzales):

wait what...

OpenStudy (anonymous):

\[SA = \pi (8)^{2} + \pi (8)(\sqrt{1433})\] \[SA = 64\pi + 8\pi(\sqrt{1433})\]

OpenStudy (anonymous):

You should use a calculator and multiply/add those values to get your surface area

OpenStudy (jessiegonzales):

200.96+8pi(37.8

OpenStudy (jessiegonzales):

200.96+25.12(37.8)

OpenStudy (anonymous):

Yes, that's the right idea. The reason you should use the pi key on your calculator is to get a closer answer: \[SA = \pi (8)^{2} + \pi (8)(\sqrt{1433}) \approx 1152.461509 \] So answer choice D. 1152 square meters